# Neural Attention Search

### v3

@@ 1 Introduction
Unlike prior works that rely on human expertise or predefined rules to identify important tokens [14, 25, 26, 29, 52, 82, 83, 93], we propose a novel approach to evaluate the importance of each token by assigning different roles to each of the tokens. For example, some tokens will be preserved until the end, while other tokens might only survive for a short amount of time. These roles measure the importance of each token and determine if they would survive within the next few tokens. Rather than pre-define a set of fixed rules for each token, we ask the model to learn this information automatically. Finding the optimal role for each token is similar to the process of neural architecture search [21, 79], where an optimal architecture is searched by the optimizer for a given task. Thereby, in this work, we jointly optimize the choice of each token and the model weights by constructing a learnable attention mask. Our approach is implicitly motivated by the one-shot neural architecture search approach [20, 53, 66], where the model parameters and architecture parameters are jointly optimized during the search process. However, our approach, Neural Attention Search (NAtS), searches for the optimal token roles jointly with the attention weights.
@@ 2.1 Attention maps for transformers
Transformers [77] computes the correlation between different tokens by mapping the input sequences into three variables, $Q$, $K$ and $V$. An attention map $A$ is first computed by pairing $Q$ and $K$ and then scaled and normalized with a softmax function. Finally, this attention map $A$ is multiplied by $V$. Additionally, a mask $M$ is attached to the Attention Map to guide the attention to focus only on certain tokens. This can be expressed as
@@
Computing the attention map of a sequence with length $\mathit{L}$ requires a complexity of $\mathcal{O}(\mathit{L}^{2})$ for both time and memory. Techniques such as KV cache lower this complexity to the $\mathcal{O}(\mathit{L}_{\text{cache}})$ during inference time, with $\mathit{L}_{\text{cache}}$ being the length of the KV cache. However, as an expense, the cost of storing the KV cache gradually becomes unnegligible with the increasing size of large language models and the context length they could handle. Therefore, lots of work has been proposed to reduce the storage and computation overheads during the inference time [50].
To reduce the computational costs of an LLM on long context information, LLMLingua [64] trains another smaller model to compress the input texts. Many other studies work on reducing the KV cache size by deciding which tokens to evict or preserve with the past attention maps. Recent works such as H2O [93], Scissorhand [56], FastGen [29], KeyFormer [6], SnapKV [52], PyramidKV [14], AadaKV [25], ChunkKV [55], and CriticalKV [26] all apply different rules to identify the important of each token and remove all the remaining tokens. Additionally, instead of simply removing the unimportant tokens, we can also merge them into the existing tokens [78, 92].
@@ 2.2 Sparse Attention
Sparse attention works on computing only a fraction of the attention map to reduce the computation costs and KV cache sizes with either pre-defined fixed patterns, including Sparse Transformers [16], Longformer [12], StreamingLLM [82], and DuoAttention [83]. Alternatively, this can be adjusted by special embedding tokens such as Longcoder [33] and SepLLM [15]. Additionally, one can also determine the types of sparse attention with a set of proxy values, including MInference [41], FlexPrefill [48], SeerAttention [28], SpargeAttention [90], and XAttention [84].
@@ 3.1 Search Space Design
Local Tokens only survive until the next Global Token appears. Therefore, models will have the full attention within each sub-sequence to summarize the local sub-sequence information into the Global Token located at the end of the sub-sequence while being sparse within the input sequence.
@@
In contrast to other causal attention maps, Figure LABEL:fig:LearnedAttention illustrates an exemplary attention mask constructed by the choices of different token types. In this case, we define the sliding window size as 4. Token 1, 4, 10 act as Global Tokens; Tokens 2, 6, 7, 8, 11 are Local Tokens; Token 3, 5, 9, 11 are Sliding Window Tokens. The Global Tokens splits the entire sequence into three subsequences that end at 4, 10, and the last index accordingly. Hence, Token 2 will only be required by Token 3 and 4. This rule applies the same for Token 6, 7, and 8, where they only interact with the tokens within the same subsequence. For the sliding window tokens, only the next three tokens could query their information, regardless of whether these tokens belong to the same sub-sequence. Only 6 out of 12 tokens are involved during inference time to make the next token prediction.
@@ 3.2 Searching for the Optimal Token Types
Specifically, we first use a linear layer (which we call Attention Score Layer) that maps each the input tensor for an attention layer $X\in\mathbb{R}^{\mathit{d}}$ to the likelihood for each option: $\alpha\in\mathbb{R}^{(\mathit{H}*\mathit{N_{\text{opts}}})}=\text{Linear}(\mathbf{x})$, where $\mathit{H}$ is the number of KV heads and $\mathit{N_{\text{opts}}}$ is the number of options in the search space. The type $\alpha$ for each token is then sampled from a Gumble-Softmax function based on the likelihood values.
@@
During the backward process, we collect the column-wise gradient values from our attention masks and apply these gradients to update the parameters from our Attention Score Layer. Different from other approaches that approximates the attention maps or attention outputs from an existing model [8, 27, 85],  NAtS can be learned directly from the target labels, and thus, can be applied for both pre-training a new sparse transformer from scratch or fine tunning an existing transformer. To control the attention map’s sparsity, we introduce a small regularization value $\lambda$ that is directly applied to the gradient for Global Token and Local Tokens to encourage more Sliding Window Tokens. We show how the backward gradients is computed and how the $\lambda$ values control the attention map sparsity in the appendix  C.
@@ 3.3 Efficient Inference with Different Token Types
During inference time, we dynamically map the input feature maps to the corresponding token types and discard the tokens no longer required by the following tokens. The Sliding Window Tokens only survive for a fixed amount of time steps. We preserve a queue in the KV cache space that only stores the most recent W tokens and mask out the non-Sliding Window Tokens: when new tokens come, we place them after the tail of the queue to replace the tokens older than W.
Similar to the vanilla transformer forward process, when new tokens arrive, we concatenate them with the existing KV caches, generating new masks and computing the attention output. After that, we start our post-update process: we first check the state of each token to decide if we can remove them or keep them in the cached values. Since different heads might disagree on the types of the same token, we record both the sizes for Global Tokens ($Size_{G}$) and Local Tokens ($Size_{L}$) for all the heads. New Sliding Window Tokens do not change these sizes since they will always be placed separately. However, when a new Global Token for any head arrives, we remove all the Local Tokens from the corresponding heads and place the new Global Token right after the existing Global Tokens and then update our $Size_{G}$ and $Size_{L}$ accordingly. The same strategy is applied when new $LocalTokens$ arrive: we place them at the end of the Local Tokens and increase the number for Local Tokens.
@@ 4.1 Training a Transformer From Scratch
We first apply NAtS to train a GPT2 small style [67] transformer model with 128M parameters from scratch on the PG-19 Language Modeling Benchmark [68] with four Nvidia H100-PCIe GPUs and evaluate them on the test sets of PG19 with a context length of 1024. Further details can be found under the appendix  E.1.
@@
We provide different hyperparameters for H2O and Streaming LLM for different sparsity (with the sliding window size of 32, 64, 128, 256 plus the same number of HH for H2O and 64, 128, 256, 512 plus 8 Sink tokens for Streaming LLM). This results in a corresponding KV budget fraction ranging from 6.25% to 50%. Meanwhile, we train NAtS with the following $\lambda$: 0,1e9,5e9,1e8,1e70,1e-9,5e-9,1e-8,1e-7.
Figure 6 shows the perplexity of the PG19 test sets from different model settings. The x-axis indicates the fraction of KV caches applied to generate the last token in the input sequence. As expected, larger $\lambda$ results in a smaller KV cache fraction. Interestingly, even if $\lambda$ is set as 0, the KV cache budget can still be reduced to around 35% with slightly higher perplexity compared to the full attention. This might indicate that the model also tends to sparsify the input information to focus more on the relevant parts from the input data [87].
@@
4.2 Fine Tune a Large Language Model
We now apply NAtS to fine-tune an existing large language model [32, 40] in the long context scenario. We construct a new training dataset that follows the construction rules of LongBench as a training set to fine-tune LLM models with NAtS. Some of LongBench’s tasks are collected from the test sets of the previous benchmarks. Hence, we first collect the training sets from these benchmarks and construct these datasets following the data structure in LongBench. Additionally, we also add the passkey-retrieval dataset introduced in DuoAttention [81]. Overall, this dataset contains 7 0007\,000 instances with a maximal context length of 16 00016\,000. Further details can be found in Appendix E.2.
@@ 4.2 Fine Tune a Large Language Model
We evaluate all the baseline results on the RULER [36] and LongBench [11]. RULER is a synthetic benchmark that evaluates a model’s long-context capabilities across 4 task categories. LongBench evaluates the model’s capacity to understand the information in different long context scenarios. Following the setting from KV Press [63], we ask all the one-time KV eviction approaches to evict the KV cache once the model receives all the context information before the question-related information arrives. For the other approaches, we evict the KV caches dynamically when new tokens are generated.
@@
The results on the Ruler Benchmark with LLama 3.1 are shown in Table  1. NAtS uses $\lambda=3e-7$ for the 25% KV size level and $\lambda=5e-8$ for the 50% KV size level. NAtS achieves nearly lossless scores across different input context lengths. When the KV budget drops to 25%, NAtS achieves 96% of the full attention scores with much smaller KV budget sizes, while the other baselines could no longer keep their performance with the decreased KV budgets. Results with Mistral-7B-Instruct-v0.3 can be found under Appendix F.
@@
Table 2: LongBench Results with 25% Budget Allocation. The Full model uses the 100% KV budgets. The numbers in the brackets for NAtS is the used value KV cache sizes. We bold the models with the best performance besides the base Full Attention Models .
For the LongBench tasks, we evaluate all the baselines with $50\%$ and $25\%$ of the KV cache sizes. We show the result with KV cache size of $25\%$ for LLama3.1-8B in Table 2. The model that is used in this task is a model with $\lambda=3e-7$. NAtS achieves the best performance on most of the datasets with (in many cases) smaller KV cache values. We provide further results, including results with Mistral-7B and  NAtS trained with other parameters, in Appendix  F.
@@ 6 Acknowledgement
Appendix A Limitations
While NAtS shows promising results that efficiently reduce the KV cache sizes and our search space already contains most of the KV cache eviction strategies. However, the construction of the attention mask is still column-wise oriented and mostly focuses on the vertical style sparse attention [41]. Hence, this work does not involve the block-sparse attention and slash attention that are widely observed in other frameworks [41, 48]. One potential future direction would be to introduce a more structured search space that allows for the construction of more sparse attention types.
Appendix B Broader Impact
Appendix C Details on Backward Propagation
C.1 Gradients for Attention Masks
$\frac{\partial g}{\partial M}$ $=e^{A}$ (9)
$\frac{\partial g}{\partial A}$ $=e^{A}\odot M$ (10)
$dg_{i:}=(diag(\frac{1}{S_{i:}})-\frac{1}{S_{i:}}P_{i:}^{T})dP_{i:}$ (11)
$dM_{i:}$ $=(diag(\frac{e^{A_{i:}}}{S_{i:}})-\frac{e^{A_{i:}}}{S_{i:}}P_{i:}^{T})dP_{i:}$ (12)
$dA$ $=dM\odot M$
$=(diag(P_{i:})-P_{i:}P^{T}_{i:})dP_{i:}$ (13)
Following FlashAttention [18], we define $D_{i}=do_{i}^{T}o_{i}$, then
$dM_{i,j}=\frac{e^{A_{i,j}}}{S_{i:}}(dP_{ij}-D_{i})$~ (14)
$dA_{i,j}=P_{i,j}(dP_{ij}-D_{i})$~ (15)
$P^{\prime}_{i,j}$ $=\min(\frac{e^{A_{i,j}}}{S_{i}},1)$ (16)
$dM_{i,j}$ $=P^{\prime}_{i,j}(dP_{ij}-D_{i})$~ (17)
$d\alpha_{i}=\sum_{j}dM_{i,j}\times M_{i,j}^{\alpha}$~ (18)
$d\alpha^{G}_{i}$ $=d\alpha^{G}_{i}-d\alpha^{G-}_{i}$ (21)
Figure  8 illustrates an example of this. Token 4 is a global token, and we search for its next Global Token (which is Token 10 in this example). Assuming that we want to change Token 4 to another role, the regions within the boundary (orange ones) are those tokens that are influenced by this swtching: given that Token 4 no longer becomes Global Token, Token 2 and 3 will be part of a larger subsequence that ranges from 1 to 10, and the attentino maps within the orange region should not be masked out. Intuitively, this gradient measures the regret that we made in order to switch one Global Token into a Local Token. A similar idea can be found in the red region. Assuming that we want to switch Token 7 to a Global Token, then Tokens 5,6 will be split into a new subsequence and we will no longer connect them with Tokens 8, 9, 10 since they belong to two different sub-sequences after the switch. Hence, this value in the red regions measures the regret if we mistakenly classify a Global Token as non-Global Tokens.
Figure 8: The gradient term $d\alpha^{G-}_{i}$ for Token 4 and 7
In practice, the $d\alpha^{G-}_{i}$ values for Global Token can be easily collected by looking at the last and next $G$ index with a scan function to collect the gradients to the corresponding positions. However, computing $d\alpha^{G-}_{i}$ with $G_{i}=0$ requires the gradient information from all the past token mask information. This might be computationally prohibitive in practice, since the valid values contained in Equation 20 are sparse, we ignore the cases for $G_{i}=0$ and only compete the dense gradeints for the case where $G_{i}=1$.
C.3 Optimizing for sparser Attention Maps
The sparse regularization term $\lambda$ is directly applied to the corresponding gradients for the Global Tokens and Local Tokens. This value should penalize the number of unmasked tokens for each row. While the number of unmasked tokens for Global Token and Local Tokens in column $i$ are $\mathit{L}-i$ and $I^{G_{\text{next}}}_{i}-i$ with $I^{G_{\text{next}}}_{i}$ defined in Section 3.2. Hence, we have:
$dG^{G_{sparse}}_{i}$ $=\lambda\times\frac{\mathit{L}-i}{\mathit{L}}$ (22)
Combining Eq.  18,  20,  22 and Eq.  23, we have:
Appendix D Integrating NAtS into FlashAttention
0: Matrices $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times d}$, token states $\mathbf{\alpha}\in\mathbb{R}^{N\times\mathit{N_{\text{opts}}}}$ , indices of the next token states $\mathbf{I^{G_{\text{next}}}}\in\mathbb{R}^{N}$ in HBM, block sizes $B_{c}$, $B_{r}$, sliding window size W
3:for $1\leq i\leq T_{r}$ do
4:   Load $\mathbf{Q}_{i}$ from HBM to on-chip SRAM.
6:  for $1\leq j\leq T_{c}$ do
8:   if Do Compute then
10:     Load $\mathbf{K}_{j},\mathbf{V}_{j}$ from HBM to on-chip SRAM.
14:   end if
15:  end for
17:  On chip, compute $L_{i}=m_{i}^{(T_{c})}+\log(\ell_{i}^{(T_{c})})$.
18:  Write $\mathbf{O}_{i}$ to HBM as the $i$-th block of $\mathbf{O}$.
19:  Write $L_{i}$ to HBM as the $i$-th block of $L$.
20:end for
21: Return the output $\mathbf{O}$ and the logsumexp $L$.
Flashattention [17] continuously loads Q, K, V values to the SRAM and only computes the attention maps within the loaded blocks.  NAtS loads the corresponding token states information $\alpha$ and construct the masks $M$ dynamically with Equation 3,  5,  4. During the backward process, Flashattention collects the row-wise gradients for computing $d\mathbf{Q}$ and column-wise gradients to compute $d\mathbf{K}$ and $d\mathbf{V}$. Hence, the gradients $d\mathbf{\alpha}$ can be computed together with $d\mathbf{K}$ and $d\mathbf{V}$. We show how  NAtS can be integrated into Flashattion2  [17] in Algorithms 1 and 2. We highlight the difference between the vanilla Flashattention2 and Flashattention2 with  NAtS in blue. Compared with vanilla Flashattention models,  NAtS adaptively skips the blocks that do not have valid mask values to avoid unnecessary computation. These invalid blocks will only be applied to update the gradients for $d\mathbf{\alpha}$.
0: Matrices $\mathbf{Q},\mathbf{K},\mathbf{V},\mathbf{O},\mathbf{dO}\in\mathbb{R}^{N\times d}$, token states $\mathbf{\alpha}\in\mathbb{R}^{\mathit{L}\times\mathit{N_{\text{opts}}}}$ , indices of the next token states $\mathbf{I^{G_{\text{next}}}}\in\mathbb{R}^{\mathit{L}}$ in HBM, vector $L\in\mathbb{R}^{N}$ in HBM, block sizes $B_{c}$, $B_{r}$, sliding window size W.
5:for $1\leq j\leq T_{c}$ do
8:  for $1\leq i\leq T_{r}$ do
13:   Check if the current block contain any valid value
14:   if Do compute then
20:   end if
22:   On chip, update $d\mathbf{\alpha}_{j}$ with Equations 24, 25, 26
23:  end for
24:  Write $\mathbf{dK}_{j},\mathbf{dV}_{j}$, and $d\mathbf{\alpha}_{j}$ to HBM.
25:end for
26: Return $\mathbf{dQ},\mathbf{dK},\mathbf{dV}$.
Appendix E Experiments Details
E.1 Detailed Hyperparameter Setting for GPT2-small Training
E.2 Collecting the Fine-tune Training Set
  • Single-Document QA: NarrativeQA [46] and Qasper [19]
  • Few-shot Leraning: TREC [51], TriviaQA [42], and SAMSum [31]
  • Code Completion: LCC [33] and RepoBench-P [54]
Appendix F Further Experimental Results
F.1 Results on Ruler Benchmark with Mistral
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
Table  3 shows the results of different optimizers on the Mistral-7B-Instruct-v0.3 model. In this case, the $\lambda$ for 25% and 50% budgets are 1e61e-6 and 3e73e-7 respectively. We found that to achieve the same sparsity level, Mistral-7B requires a larger sparse regularization value $\lambda$ compared to the Llama 3.1 model. This might indicate that the attention map values in Mistral are more evenly distributed, and thus we need a larger $\lambda$ to force more tokens to only focus on the local information. The maximal context length of Mistral 7B is $32k$. We only evaluate the models with a maximal length of $32k$. The result is consistent with the results shown in Table 1.  NAtS consistently outperforms the other baselines with smaller budgets.
F.2 Further results on LongBench
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
Table 4: LongBench Results with 25% Budget Size on Mistral-7B. .
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
Table 5: LongBench Results with 50% Budget Size on LLama 3.1 8B .
Full Duo SLLM H2O Snap Ada Chunk Critical Pyradmid MoA NAtS
Table 6: LongBench Results with 50% Budget Size on Mistral-7B. .
Appendix G Ablation Study
G.1 Sparse Regularization Term  $\lambda$
Full NAtS 1e-6 NAtS 5e-7 NAtS 3e-7 NAtS 1e-7 NAtS 5e-8
Qasper 24.7 31.39(15%) 28.6(21%) 23.76(27%) 25.17(45%) 26.97(58%)
Musique 11.6 10.78(9%) 9.87(13%) 11.42(18%) 11.89(36%) 11.31(48%)
QMSum 23.3 23.37(8%) 23.09(12%) 23.11(17%) 22.6(33%) 22.74(45%)
TREC 72.5 68.0(14%) 70.5(20%) 72.0(26%) 73.5(44%) 73.0(56%)
SAMSum 43.7 43.98(8%) 44.0(12%) 44.21(16%) 42.61(30%) 43.84(41%)
LSHT 46.5 46.0(8%) 44.0(12%) 47.5(16%) 45.5(31%) 45.5(43%)
LCC 54.1 55.46(24%) 54.53(31%) 53.54(38%) 53.27(57%) 54.4(69%)
G.2 Sliding Window Length W
Full NAtS 64 NAtS 128 NAtS 256 NAtS 512
NarrativeQA 31.4 21.61(33%) 30.76(15%) 30.53(13%) 28.96(12%)
Qasper 24.7 17.92(34%) 24.1(28%) 23.76(27%) 28.24(29%)
MultiFieldQA-en 29.5 19.37(34%) 29.02(26%) 28.94(24%) 25.87(25%)
MultiFieldQA-zh 60.0 35.49(35%) 60.37(22%) 61.18(22%) 58.08(26%)
HotpotQA 17.1 13.99(33%) 16.41(23%) 16.06(20%) 17.44(18%)
2WikiQA 16.6 11.52(34%) 16.69(26%) 17.05(23%) 16.48(24%)
Musique 11.6 9.42(33%) 11.62(22%) 11.42(18%) 10.89(16%)
DuReader (zh) 35.6 22.26(34%) 35.75(19%) 34.28(17%) 33.91(16%)
GovReport 34.3 26.63(34%) 33.89(22%) 33.82(19%) 33.05(19%)
QMSum 23.3 20.46(33%) 23.52(19%) 23.11(17%) 22.66(16%)
MultiNews 27.1 23.78(36%) 26.73(33%) 26.72(33%) 26.46(40%)
VCSUM (zh) 16.4 13.8(34%) 16.1(15%) 15.61(14%) 15.9(15%)
TREC 72.5 46.5(34%) 73.0(31%) 72.0(26%) 70.5(25%)
TriviaQA 91.2 84.44(34%) 91.89(26%) 91.61(22%) 91.66(21%)
SAMSum 43.7 38.97(34%) 43.62(18%) 44.21(16%) 44.47(17%)
LSHT 46.5 26.5(34%) 47.0(18%) 47.5(16%) 45.5(15%)
Passage Count 6.6 0.0(34%) 5.75(23%) 7.12(20%) 5.66(18%)
LCC 54.1 39.13(36%) 54.92(39%) 53.54(38%) 54.15(44%)
RepoBench-P 51.4 36.87(34%) 51.89(32%) 53.48(28%) 51.34(26%)
Table 8: Ablation Study of Sliding Window Length W for Llama3
The result is shown in Table 8. Overall, all the approaches perform similarly. However, a smaller sliding window size generally results in an overall larger KV cache size. A reduced sliding window size would force the model to apply more Global Tokens and Local Tokens to construct the mid-range correlation since this distance cannot be covered by the sliding window tokens. However, as the number of sliding window size further increases, this compression rate might saturate since the remaining tokens might always require a long-range correlation whose distance is much larger than the sliding window size (e.g., these tokens might require the correlation between two tokens whose distances are larger than 1k or even more).
Appendix H KV size Distributions
(a) KV size of LLama on Narrative QA Dataset
(b) KV size of LLama on Multi-News Dataset
Figrues  10(a) and  10(b) show the KV cache size distributions for Llama 3.1 8B on the Narrative QA and Multi News Datasets. The models tend to preserve more KV caches in the shallower layers and ignore the KV caches in the intermediate layers. However, the model still preserves several KV caches near the output layers. Additionally, even in the same layer, the distributions of the KV cache sizes are not evenly distributed in each layer. Some heads are always preferred while the others might be dropped as the context or $\lambda$ changes.
(c) KV size of Mistral on NarrativeQA Dataset
(d) KV size of Mistral on Multi-News Dataset
Figures  10(c) and  10(d) show the KV cache size distributions from the Mistral model. Although the two models share some similar trends, e.g., both models tend to preserve more tokens in the shallower and the last few layers. However, compared to Llama, the KV caches for the Mistral model are more unevenly distributed and tend to gather towards some specific heads. This highlights the importance of adapting different KV eviction strategies to different models.

### v4

@@ 1 Introduction
Unlike prior works that rely on human expertise or predefined rules to identify important tokens [15, 27, 28, 31, 54, 84, 85, 96], we propose a novel approach to evaluate the importance of each token by assigning different roles to each of the tokens. For example, some tokens will be preserved until the end, while other tokens might only survive for a short amount of time. These roles measure the importance of each token and determine if it would survive within the next few tokens. Rather than pre-defining a set of fixed rules for each token, we ask the model to learn this information automatically. Finding the optimal role for each token is similar to the process of neural architecture search [23, 81], where an optimal architecture is searched by the optimizer for a given task. Thereby, in this work, we jointly optimize the choice of each token and the model weights by constructing a learnable attention mask. Our approach is implicitly motivated by the one-shot neural architecture search approach [22, 55, 68], where the model parameters and architecture parameters are jointly optimized during the search process. However, our approach, Neural Attention Search (NAtS), searches for the optimal token roles jointly with the attention weights.
@@ 2.1 Attention maps for transformers
Transformers [79] computes the correlation between different tokens by mapping the input sequences into three variables, $Q$, $K$, and $V$. An attention map $A$ is first computed by pairing $Q$ and $K$ and then scaled and normalized with a softmax function. Finally, this attention map $A$ is multiplied by $V$. Additionally, a mask $M$ is attached to the Attention Map to guide the attention to focus only on certain tokens. This can be expressed as
@@
Computing the attention map of a sequence with length $\mathit{L}$ requires a complexity of $\mathcal{O}(\mathit{L}^{2})$ for both time and memory. This complexity can be lowered to $\mathcal{O}(\mathit{L}_{\text{cache}})$ during inference time, with $\mathit{L}_{\text{cache}}$ being the length of the KV cache. However, as an expense, the cost of storing the KV cache gradually becomes unnegligible with the increasing size of large language models and the context length they could handle. Therefore, lots of work has been proposed to reduce the storage and computation overheads during the inference time [52].
To reduce the computational costs of an LLM on long context information, LLMLingua [66] trains another smaller model to compress the input texts. Many other studies work on reducing the KV cache size by deciding which tokens to evict or preserve with the past attention maps. Recent works such as H2O [96], Scissorhand [58], FastGen [31], KeyFormer [6], SnapKV [54], PyramidKV [15], AadaKV [27], ChunkKV [57], and CriticalKV [28] all apply different rules to identify the importance of each token and remove all the remaining tokens. Additionally, instead of simply removing the unimportant tokens, we can also merge them into the existing tokens [80, 95].
@@ 2.2 Sparse Attention
Sparse attention works on computing only a fraction of the attention map to reduce the computation costs and KV cache sizes with either pre-defined fixed patterns, including Sparse Transformers [17], Longformer [13], StreamingLLM [84], and DuoAttention [85]. Alternatively, this can be adjusted by special embedding tokens such as Longcoder [35] and SepLLM [16]. Additionally, one can also determine the types of sparse attention with a set of proxy values, including MInference [43], FlexPrefill [50], SeerAttention [30], SpargeAttention [93], and XAttention [86]. However, these approaches aim to approximate the attention output of each layer from a pre-trained dense transformer that only involves the QK matrices. In contrast, NAtS learns the token roles directly from the final loss function and does not necessarily need to approximate the existing attention output. Hence, NAtS can be applied to both training from scratch or approximating the outputs from an existing model. Additionally, the existing approaches can only be applied to approximating the attention maps during the pre-filling stages and still need to preserve the entire KV cache during the decoding stage, making them less profitable for reasoning models where the model needs to generate lots of tokens during the decoding stage [21]. In contrast, NAtS only checks the roles for each token and does not necessarily need to link to the actual attention map values. Hence, it can be applied to both pre-filling and decoding stages while keeping the KV cache sizes to a low level.
@@ 3.1 Search Space Design
Local Tokens only survive until the next Global Token appears. Therefore, models will have the full attention within each sub-sequence to summarize the local sub-sequence information into the Global Token which is located at the end of the sub-sequence. Meanwhile, the model will be sparse within the input sequence. Local Tokens are considered as tokens that only provide lower-level information that helps the model to understand the local context information, but might not help further outside this context. This provides the model with a flexible interface to control its sparsity based on the input context information.
@@
In contrast to other causal attention maps, Figure LABEL:fig:LearnedAttention illustrates an exemplary attention mask constructed by the choices of different token types. In this case, we define the sliding window size as 4. Token 1, 4, 10 act as Global Tokens; Tokens 2, 6, 7, 8, 11 are Local Tokens; Token 3, 5, 9, 11 are Sliding Window Tokens. The Global Tokens splits the entire sequence into three subsequences that end at 4, 10, and the last index, respectively. Hence, Token 2 will only be required by Token 3 and 4. This rule applies the same for Token 6, 7, and 8, where they only interact with the tokens within the same subsequence. For the sliding window tokens, only the next three tokens could query their information, regardless of whether these tokens belong to the same sub-sequence. Only 6 out of 12 tokens are involved during inference time to make the next token prediction.
@@ 3.2 Searching for the Optimal Token Types
Specifically, we first use a linear layer (which we call Attention Score Layer) that maps each the input tensor for an attention layer $X\in\mathbb{R}^{\mathit{d}}$ to the likelihood for each option: $\alpha\in\mathbb{R}^{(\mathit{H}*\mathit{N_{\text{opts}}})}=\text{Linear}(\mathbf{x})$, where $\mathit{H}$ is the number of KV heads and $\mathit{N_{\text{opts}}}$ is the number of options in the search space. The type $\alpha$ for each token is then sampled from a Gumble-Softmax function based on the likelihood values. Given that the number of options $\mathit{N_{\text{opts}}}$ is normally much smaller than the model head dimensions, this linear layer only introduces minimal overhead compared to the QKV linear layers.
@@
During the backward process, we collect the column-wise gradient values from our attention masks and apply these gradients to update the parameters from our Attention Score Layer. Different from other approaches that approximate the attention maps or attention outputs from an existing model [9, 29, 87],  NAtS can be learned directly from the target labels, and thus, can be applied for both pre-training a new sparse transformer from scratch or fine-tuning an existing transformer. To control the attention map’s sparsity, we introduce a small regularization value $\lambda$ that is directly applied to the gradient for Global Token and Local Tokens to encourage more Sliding Window Tokens. This regularization value $\lambda$ can be considered as a value similar to a soft threshold, pushing the tokens whose contributions in the attention maps outside the scope of the sliding window sizes with lower levels towards Sliding Window Tokens. Hence, we could efficiently reduce the corresponding overall computational costs.
We show how the backward gradients are computed and how the $\lambda$ values control the attention map sparsity in the appendix  C.
@@ 3.3 Efficient Inference with Different Token Types
NAtS can be applied to both pre-filling and decoding stages. During the pre-filling stage, we first compute the attention values for all the global and local tokens. We then compute attention map values for the sliding window tokens separately to fully utilize the parallelism of GPU architectures.
During the decoding stage, we dynamically map the input feature maps to the corresponding token types and discard the tokens no longer required by the following tokens once a new token arrives. The Sliding Window Tokens only survive for a fixed number of time steps. We preserve a queue in the KV cache space that only stores the most recent W tokens and mask out the non-Sliding Window Tokens: when new tokens come, we place them after the tail of the queue to replace the tokens older than W.
Similar to the vanilla transformer forward process, when new tokens arrive, we concatenate them with the existing KV caches, generating new masks and computing the attention output. After that, we start our post-update process: we first check the state of each token to decide if we can remove them or keep them in the cached values. Since different heads might disagree on the types of the same token, we record both the sizes for Global Tokens ($Size_{G}$) and Local Tokens ($Size_{L}$) for all the heads. New Sliding Window Tokens do not change these sizes since they will always be placed separately. However, when a new Global Token for any head arrives, we remove all the Local Tokens from the corresponding heads and place the new Global Token right after the existing Global Tokens and then update our $Size_{G}$ and $Size_{L}$ accordingly. The same strategy is applied when new $LocalTokens$ arrive: we place them at the end of the Local Tokens and increase the number for Local Tokens. A detailed updating process can be found under Appendix E.
@@ 4.1 Training a Transformer From Scratch
We first apply NAtS to train a GPT2 small style [69] transformer model with 128M parameters from scratch on the PG-19 Language Modeling Benchmark [70] with four Nvidia H100-PCIe GPUs and evaluate it on the test sets of PG19 with a context length of 1024. Further details can be found under the appendix  F.1.
@@
We provide different hyperparameters for H2O and Streaming LLM for different sparsity (with the sliding window size of 32, 64, 128, 256, plus the same number of HH for H2O and 64, 128, 256, 512, plus 8 Sink tokens for Streaming LLM). This results in a corresponding KV budget fraction ranging from 6.25% to 50%. Meanwhile, we train NAtS with the following $\lambda$: 0,1e9,5e9,1e8,1e70,1e-9,5e-9,1e-8,1e-7. We run each experiment with three different seeds.
Figure 6 shows the perplexity and the corresponding standard deviation (as error bars) of the PG19 test sets from different model settings across different seeds. The x-axis indicates the fraction of KV caches applied to generate the last token in the input sequence. As expected, a larger $\lambda$ results in a smaller KV cache fraction, while the performance is relatively stable across different seeds. Interestingly, even if $\lambda$ is set as 0, the KV cache budget can still be reduced to around 35% with slightly lower perplexity compared to the full attention. This might indicate that the model also tends to sparsify the input information to focus more on the relevant parts from the input data [89].
@@
4.2 Fine-Tune a Large Language Model
We now apply NAtS to fine-tune an existing large language model [34, 42] in the long context scenario. We construct a new training dataset that follows the construction rules of LongBench as a training set to fine-tune LLM models with NAtS. Some of LongBench’s tasks are collected from the test sets of the previous benchmarks. Hence, we first collect the training sets from these benchmarks and construct these datasets following the data structure in LongBench. Additionally, we also add the passkey-retrieval dataset introduced in DuoAttention [83]. Overall, this dataset contains 7 0007\,000 instances with a maximal context length of 16 00016\,000. Further details can be found in Appendix F.2.
@@ 4.2 Fine-Tune a Large Language Model
We evaluate all the baseline results on the RULER [38] and LongBench [12]. RULER is a synthetic benchmark that evaluates a model’s long-context capabilities across 4 task categories. LongBench evaluates the model’s capacity to understand information in different long context scenarios. Following the setting from KV Press [65], we ask all the one-time KV eviction approaches to evict the KV cache once the model receives all the context information before the question-related information arrives. For the other approaches, we evict the KV caches dynamically when new tokens are generated.
@@
The results on the Ruler Benchmark with LLama 3.1 are shown in Table 1. NAtS uses $\lambda=3e-7$ for the 25% KV size level and $\lambda=5e-8$ for the 50% KV size level. NAtS achieves nearly lossless scores across different input context lengths. When the KV budget drops to 25%, NAtS achieves 96% of the full attention scores with much smaller KV budget sizes, while the other baselines could no longer keep their performance with the decreased KV budgets. Results with Mistral-7B-Instruct-v0.3 can be found under Appendix G.
@@
Table 2: LongBench Results with 25% Budget Allocation. The Full model uses the 100% KV budgets. The numbers in the brackets for NAtS are the used value KV cache sizes. We bold the models with the best performance besides the base Full Attention Models .
For the LongBench tasks, we evaluate all the baselines with $50\%$ and $25\%$ of the KV cache sizes. We show the result with KV cache size of $25\%$ for LLama3.1-8B in Table 2. The model that is used in this task is a model with $\lambda=3e-7$. NAtS achieves the best performance on most of the datasets with (in many cases) smaller KV cache values. We provide further results, including results with Mistral-7B and  NAtS trained with other parameters, in Appendix  G.